# A Momentum Balance Correction to the Non-Conservative One-Fluid Formulation in Boiling Flows using Volume-of-Fluid

### v1

@@ 5 Coupling with Energy Equation
This section addresses the fully-coupled system involving the energy equation and the calculation of the interfacial mass flux. Section 5.1 looks at the growth of a static vapor bubble in superheated liquid, while Section 5.2 considers a dynamic setup where a bubble rises in superheated liquid due to the effects of buoyancy. Then, Section 5.3 examines a two-dimensional film boiling configuration. These analyses look at the correct solution of the energy equation and the effects of the momentum imbalance and its corrections. Also, they involve a variety of fluid properties and show that the flow solver behaves well at, e.g., low and high density ratios and different surface tension coefficients. The relevant properties of the fluids considered here are listed in Table 3.
@@ 5.1 Static Bubble Growth in Superheated Liquid
Similar to other works [13, 17], a fluid with the properties of fluid A shown in Table 3 is considered, which is defined for numerical testing purposes and may not resemble any real fluid. The bubble is immersed in a superheated liquid with $T_{\infty}=3$ K. These conditions may arise when a sudden decrease in system pressure causes the boiling of a previously saturated liquid at a higher pressure. The assumption that the change in $T_{sat}$ is small holds, and the effect of temperature on the fluid properties is neglected, i.e., remain constant.
@@
The evolution of the bubble radius, temperature radial profiles at $t=1$ s and $t=2$ s (extracted along the $z$ direction), and average mass flux are shown in Figure 11. The analytical solution is captured and converging trends are observed. Still, the bubble radius is directly related to the volume and the limitations of the plane shift strategy shown in Section 4.2 may resurface if very high accuracy in the calculation of $\dot{m}^{\prime}$ is achieved. Table 5 reports convergence rates of the bubble radius and average mass flux between first and second order, which are in line with previous works [13, 17].
@@
Table 5: Relative errors, $|\phi_{n}-\phi_{e}|/\phi_{e}$, and convergence rate of the bubble radius and average mass flux at $t=2$ s. The analytical radius and mass flux are, respectively, $R_{e}=0.234081$ m and $\dot{m}^{\prime}_{e}=0.0150569$ kg/(m2s).
In this specific test, the error introduced by the jump in viscous stresses is significant given the relatively high viscosity of fluid A. In other words, the pressure jump and pressure field are inconsistent with the proposed modeling neglecting corrections to the viscous jump, and a pressure spike is observed across the interface. Thus, the analytical solution, similar to the one presented in Section 4.2, cannot be recovered. Nonetheless, a consistent velocity field, i.e., Stefan flow, is still obtained due to the static configuration. The bubble remains centered in the computational domain, as indicated by the results presented in Figure 11. In Section 5.3, however, it is shown that the inconsistent viscous jump is critical in a dynamic setup, emphasizing the need to develop consistent momentum balance frameworks in future works.
@@ 5.2 Rising Bubble in Superheated Liquid
The bubble growth is measured in terms of the average radius from the bubble diameters in each direction computed as $R_{b}=(d_{x}+d_{y}+2d_{z})/8$. Figure 12 reports the normalized radius $R^{*}=R_{b}/(2\beta\sqrt{\alpha_{L}})$ where $\beta=5.39869$ is the growth constant from Scriven’s solution, i.e., Eq. (42), and $\alpha_{L}=k_{L}/(\rho_{L}c_{p,L})$ the liquid’s thermal diffusivity. Results for different mesh resolutions are shown and are compared to available data from numerical [8, 42] and experimental [57] works, as well as Scriven’s solution for bubble growth under zero gravity. The sensitivity to mesh resolution is evident. That is, if the thermal boundary layer is under-resolved, the evaluation of $\dot{m}^{\prime}$ affects the bubble growth and the overall dynamics of the problem. For the mesh with a grid spacing of $\Delta x=25$ $\mu$m, equivalent to an initial bubble resolution of 16 c/d, the evaporation of the bubble is significantly underestimated. In contrast, a resolution of $\Delta x=12.5$ $\mu$m (i.e., initially 32 c/d) is more in line with previous numerical studies and follows the reference data better. Due to limitations in the available computational resources, finer mesh resolutions have not been considered. Thus, certain parameters shown in this section, such as $R^{*}$, are not guaranteed to have converged, probably requiring a mesh resolution closer to $\Delta x=7$  $\mu$m based on previous works [8, 42].
@@
In Figure 12, substantial differences can be seen in the evolution of $R^{*}$ between three different works with very similar resolutions, namely the recent results using a grid spacing of 12.5 $\mu$m and two previous works using a resolution of 15.6 $\mu$m [8, 42]. While these differences can partially be justified by the thermal boundary layer resolution and the accuracy in evaluating $\dot{m}^{\prime}$ in each work, uncertainties remain that may be linked to the different code frameworks. This highlights the intrinsic difficulties that arise in two-phase solvers once phase change is involved. Figure 13 presents the evolution of $(d_{x}+d_{y})/2$, $d_{z}$ and the aspect ratio $2d_{z}/(d_{x}+d_{y})$ obtained with the resolution of 12.5 $\mu$m. After a brief growth stage where the bubble diameter increases almost uniformly in all directions, it begins to grow faster in the radial direction outward of the $z$ axis, evolving into a flat ellipsoid or disk (see Figure 14).
@@
The impact of mesh resolution on the evolution of the bubble’s shape is also visualized in Figure 14, which compares side by side the bubble location and shape obtained with the resolutions of 25 $\mu$m and 12.5 $\mu$m. Similar to the discussion around Figure 12, the 12.5 $\mu$m mesh resolves the thermal boundary layer better, enhancing the calculation of $\dot{m}^{\prime}$ and increasing the bubble volume faster than with 25 $\mu$m. For reference, Figure 15 depicts the contours on the $yz$ plane at $x=0$ of temperature, mass flux, velocity magnitude and pressure around the bubble at $t=45.6$ ms obtained with the grid spacing of 12.5 $\mu$m. The thin thermal boundary layer along the bubble’s upper side contrasts sharply with the bubble size. The vaporization mass flux is stronger there, reaching a peak value around 0.02 kg/(m2s) near the edge of the bubble. Instead, almost no vaporization occurs at the bottom of the bubble due to the low temperature region in the tail. Indicated by the lower pressure and the velocity magnitude contours, a recirculation region is identified along the bubble’s edge as in previous works [8].
@@
The shift in rise velocity is caused by the momentum balance corrections, as shown in Figure 16(b). Following the previous discussion on the grid sensitivity of $w_{b}$, the coarse mesh is used to analyze the early $w_{b}$ trends. Three cases are shown together with the reference data [8]: (a) with all corrections, (b) only adding $\mathbf{f}_{NC}$ and $\mathbf{f}_{\dot{m}^{\prime}}$, and (c) with no corrections. The latter is more aligned with the numerical approach from [8] to solve the one-fluid Navier-Stokes equations and reproduces more closely the $w_{b}$ trends shown in their work. As the momentum balance corrections are progressively considered, first by only adding $\mathbf{f}_{NC}$ and $\mathbf{f}_{\dot{m}^{\prime}}$ and then by also adding the first predictor step to shift the Stefan flow, an increase in $w_{b}$ occurs. In any case, all the reported velocities fall within expected values [57, 58]. This result is somewhat surprising given that the maximum value of $\dot{m}^{\prime}$ is around 0.02 kg/(m2s) throughout the simulation, leading to believe that the momentum balance corrections are negligible, i.e., $(\dot{m}^{\prime})^{2}(\rho_{G}^{-1}-\rho_{L}^{-1})$ is of $\mathcal{O}(10^{-4})$ Pa. However, they still impact the pressure field solution and the flow dynamics.
Additional insights can be gained by looking at various details of the pressure field. Figure 17 shows the pressure contours at $t=7.6$ ms on a $yz$ plane at $x=0$, focusing on the pressure around the bubble and considering the three momentum balance modeling frameworks shown in Figure 16(b). In this snapshot, the bubble size and position is almost identical for all frameworks. Nonetheless, different pressure distributions and magnitudes are observed, with the simulation without momentum balance corrections seemingly predicting a lower pressure in the liquid overall. This is corroborated in Figure 18, which shows the pressure extracted along the diagonal of an $xy$ plane located at $z=0.00625$ mm, i.e., from $\mathbf{x}_{i}=(0,0,0.00625)$ mm to $\mathbf{x}_{f}=(4,4,0.00625)$ mm, at the same time $t$. Under the assumption that the bubble and its motion have negligible influence on the pressure at the considered depth, the hydrostatic pressure given the liquid column is 118.772 Pa. Indeed, the omission of the momentum balance corrections induces a lower pressure, while the addition of $\mathbf{f}_{NC}$, $\mathbf{f}_{\dot{m}^{\prime}}$ and the Stefan flow shift improves the agreement.
Figure 17: Pressure field on the $yz$ plane with $x=0$ at $t=7.6$ ms of the rising ethanol bubble obtained with a grid spacing of 12.5 $\mu$m. The results without momentum balance corrections (left), only with the body forces $\mathbf{f}_{NC}$ and $\mathbf{f}_{\dot{m}^{\prime}}$ (center), and with all corrections (right) are shown. The interface location is visualized by the iso-contour with $C=0.5$.
@@
However, further inspection reveals that the pressure field is affected by considerable oscillations if no momentum balance corrections are implemented, similar to Figure 6(c). Figure 19 shows the average pressure inside the bubble (i.e., from all pure gas cells), the average pressure around the bubble (i.e., from all pure liquid cells directly neighboring the interface) and the average pressure jump calculated by subtracting the two previous quantities. After an initial transient period due to the relaxation of the initial conditions, the pressure inside and around the bubble obtained with all corrections implemented behaves smoothly over time. In contrast, frequent pressure oscillations are observed if the momentum balance is affected by the numerical approach. Here, the addition of only the body forces $\mathbf{f}_{NC}$ and $\mathbf{f}_{\dot{m}^{\prime}}$ helps reducing the pressure oscillations although nearly complete mitigation is only possible once the Stefan flow shift is considered. It can be observed that at $t=7.6$ ms, the solution without corrections is predicting a smaller pressure than the trend, i.e., the observations from Figures 17 and 18. Moreover, the imbalance oscillations translate into pressure jump oscillations. Notably, taking proper care of the momentum balance is reflected in the pressure jump during the initial transient where, despite the pressure field oscillating, the pressure jump trend is relatively smooth.
@@
Lastly, it is worth noting that the trends of the pressure field quantities reported in Figure 19 are very similar during the early stages despite the oscillations introduced by the formulation without momentum balance corrections. This is more aligned with the expectation that the corrections are, in terms of magnitude, negligible for this problem. However, different rise velocities are reported. A possible explanation is that the skewness of these oscillations, which will depend on the evolving shape and how the interface intersects with the mesh, is important. Then, the cumulative evolution results in different rise velocities. This observation merits further analysis.
5.3 Two-Dimensional Film Boiling
Two-dimensional film boiling is considered as an additional benchmark [59, 60, 17]. Another test fluid with the properties of fluid B shown in Table 3 is used. Following Rayleigh-Taylor instability theory, the vapor film is initialized with a perturbation of wavelength $\lambda_{d}$ to trigger the most unstable mode. The evaporation of the liquid increases the vapor volume and the dynamics are dominated by the combined effect of the momentum balance (e.g., surface tension) and buoyancy, forming a bubble that rises due to gravity (i.e., $\mathbf{g}=(0,-9.81)$ m/s2).
The computational domain is a box of size $[0,\lambda_{d}/2]\times[0,\lambda_{d}]$ with symmetric boundaries at $x=0$ and $x=\lambda_{d}/2$, a wall with no-slip boundary conditions at $y=0$ and outflow boundary conditions at $y=\lambda_{d}$. For the pressure solver, homogeneous Neumann boundary conditions are imposed at the bottom and top boundaries with a reference pressure of 0 Pa at $\mathbf{x}_{p}=(0,\lambda_{d})$. For this problem, PLIC is used since no major differences are observed in the curvature computation and interface evolution with respect to PPIC; thus, favoring computational savings. The liquid remains at $T_{sat}$ and the wall temperature is fixed at $T_{wall}=T_{sat}+5$ K. The vapor film, which sits on top of the wall, is initialized with a linear temperature profile, $T_{0}$, between the initial interface location, $y_{0}$, and the wall. $\lambda_{d}$, $T_{0}$ and $y_{0}$ are given, respectively, by Eqs. (44), (45) and (46) (see [17] for more details). For the given fluid properties, $\lambda_{d}=78.6844$ mm.

### v2

@@ 5 Coupling with Energy Equation
This section addresses the fully-coupled system involving the energy equation and the calculation of the interfacial mass flux. Section 5.1 looks at the growth of a static vapor bubble in superheated liquid, while Section 5.2 considers a dynamic setup where a bubble rises in superheated liquid due to the effects of buoyancy. These analyses look at the correct solution of the energy equation and the effects of the momentum imbalance and its corrections. Also, they involve a variety of fluid properties and show that the flow solver behaves well at, e.g., low and high density ratios and different surface tension coefficients. The relevant properties of the fluids considered here are listed in Table 3.
@@ 5.1 Static Bubble Growth in Superheated Liquid
Similar to other works [13, 17], a fluid with the properties of fluid A shown in Table 3 is considered, which is defined for numerical testing purposes and may not resemble any real fluid. The bubble is immersed in a superheated liquid with $T_{\infty}=3$ K. These conditions may arise when a sudden decrease in system pressure causes the boiling of a previously saturated liquid at a higher pressure. The assumption that the change in $T_{sat}$ is small holds, and the effect of temperature on the fluid properties is neglected, i.e., remain constant. In terms of the Jakob number, $Ja=[\rho_{L}c_{p,L}(T_{\infty}-T_{sat})]/\rho_{G}h_{LV}$, this configuration with the properties of fluid A corresponds to $Ja=0.5$.
@@
The evolution of the bubble radius, temperature radial profiles at $t=1$ s and $t=2$ s (extracted along the $z$ direction), and average mass flux are shown in Figure 12. The analytical solution is captured and converging trends are observed. Table 5 reports convergence rates of the bubble radius and average mass flux between first and second order, which are in line with previous works [13, 17]. Still, the bubble radius is directly related to the volume and the limitations of the plane shift strategy shown in Section 4.2 (i.e., low volume convergence rates with fixed mass flux) may resurface once the mass flux errors become very small. That is, the volume error clearly decreases with mesh refinement as long as the calculated $\dot{m}^{\prime\prime}$ improves significantly.
@@
Table 5: Relative errors, $|\phi_{n}-\phi_{e}|/\phi_{e}$, and convergence rate of the bubble radius and average mass flux at $t=2$ s. The analytical radius and mass flux are, respectively, $R_{e}=0.234081$ m and $\dot{m}^{\prime\prime}_{e}=0.0150569$ kg/(m2s).
When accounting for viscosity, the error introduced by the jump in viscous stresses is significant given the relatively high viscosity of fluid A. If viscosity is neglected as in the original work by Scriven [58], the correct pressure jump is obtained. In other words, the pressure jump and pressure field are inconsistent with the proposed modeling neglecting corrections to the viscous jump, and a pressure spike is observed across the interface. Moreover, the pressure solution suffers from the same limitations discussed in Section 4.2 regarding the quasi-steady interface behavior assumption. Nonetheless, a consistent velocity field, i.e., Stefan flow, is still obtained due to the static configuration. The bubble remains centered in the computational domain, as indicated by the results presented in Figure 12. In Section 6, however, it is shown that the inconsistent viscous jump is critical in a dynamic setup, emphasizing the need to develop consistent momentum balance frameworks in future works.
@@ 5.2 Rising Bubble in Superheated Liquid
The bubble growth is measured in terms of the average radius from the bubble diameters in each direction computed as $R_{b}=(d_{x}+d_{y}+2d_{z})/8$. Figure 13 reports the normalized radius $R^{*}=R_{b}/(2\beta\sqrt{\alpha_{L}})$ where $\beta=5.39869$ is the growth constant from Scriven’s solution, i.e., Eq. (42), and $\alpha_{L}=k_{L}/(\rho_{L}c_{p,L})$ the liquid’s thermal diffusivity. Results for different mesh resolutions are shown and are compared to available data from numerical [8, 42] and experimental [59] works, as well as Scriven’s solution for bubble growth under zero gravity. The sensitivity to mesh resolution is evident. That is, if the thermal boundary layer is under-resolved, the evaluation of $\dot{m}^{\prime\prime}$ affects the bubble growth and the overall dynamics of the problem. For the mesh with a grid spacing of $\Delta x=25$ $\mu$m, equivalent to an initial bubble resolution of 16 c/d, the evaporation of the bubble is significantly underestimated. In contrast, a resolution of $\Delta x=12.5$ $\mu$m (i.e., initially 32 c/d) is more in line with previous numerical studies and follows the reference data better. Due to limitations in the available computational resources, finer mesh resolutions have not been considered. Thus, certain parameters shown in this section, such as $R^{*}$, are not guaranteed to have converged, probably requiring a mesh resolution closer to $\Delta x=7$  $\mu$m based on previous works [8, 42].
@@
In Figure 13, substantial differences can be seen in the evolution of $R^{*}$ between three different works with very similar resolutions, namely the recent results using a grid spacing of 12.5 $\mu$m and two previous works using a resolution of 15.6 $\mu$m [8, 42]. While these differences can partially be justified by the thermal boundary layer resolution and the accuracy in evaluating $\dot{m}^{\prime\prime}$ in each work, uncertainties remain that may be linked to the different code frameworks. This highlights the intrinsic difficulties that arise in two-phase solvers once phase change is involved. Figure 14 presents the evolution of $(d_{x}+d_{y})/2$, $d_{z}$ and the aspect ratio $2d_{z}/(d_{x}+d_{y})$ obtained with the resolution of 12.5 $\mu$m. After a brief growth stage where the bubble diameter increases almost uniformly in all directions, it begins to grow faster in the radial direction outward of the $z$ axis, evolving into a flat ellipsoid or disk (see Figure 15).
@@
The impact of mesh resolution on the evolution of the bubble’s shape is also visualized in Figure 15, which compares side by side the bubble location and shape obtained with the resolutions of 25 $\mu$m and 12.5 $\mu$m. Similar to the discussion around Figure 13, the 12.5 $\mu$m mesh resolves the thermal boundary layer better, enhancing the calculation of $\dot{m}^{\prime\prime}$ and increasing the bubble volume faster than with 25 $\mu$m. For reference, Figure 16 depicts the contours on the $yz$ plane at $x=0$ of temperature, mass flux, velocity magnitude and pressure around the bubble at $t=45.6$ ms obtained with the grid spacing of 12.5 $\mu$m. The thin thermal boundary layer along the bubble’s upper side contrasts sharply with the bubble size. The vaporization mass flux is stronger there, reaching a peak value around 0.02 kg/(m2s) near the edge of the bubble. Instead, almost no vaporization occurs at the bottom of the bubble due to the low temperature region in the tail. Indicated by the lower pressure and the velocity magnitude contours, a recirculation region is identified along the bubble’s edge as in previous works [8].
@@
The shift in rise velocity is caused by the momentum balance corrections, as shown in Figure 17(b). Following the previous discussion on the grid sensitivity of $w_{b}$, the coarse mesh is used to analyze the early $w_{b}$ trends. Three cases are shown together with the reference data [8]: (a) with all corrections, (b) only adding $\mathbf{f}_{NC}$ and $\mathbf{f}_{\dot{m}^{\prime\prime}}$, and (c) with no corrections. The latter is more aligned with the numerical approach from [8] to solve the one-fluid Navier-Stokes equations and reproduces more closely the $w_{b}$ trends shown in their work. As the momentum balance corrections are progressively considered, first by only adding $\mathbf{f}_{NC}$ and $\mathbf{f}_{\dot{m}^{\prime\prime}}$ and then by also adding the first predictor step to shift the Stefan flow, an increase in $w_{b}$ occurs. In any case, all the reported velocities fall within expected values [59, 60]. This result is somewhat surprising given that the maximum value of $\dot{m}^{\prime\prime}$ is around 0.02 kg/(m2s) throughout the simulation, leading to believe that the momentum balance corrections are negligible, i.e., $(\dot{m}^{\prime\prime})^{2}(\rho_{G}^{-1}-\rho_{L}^{-1})$ is of $\mathcal{O}(10^{-4})$ Pa. However, they still impact the pressure field around the bubble and the flow dynamics.
Additional insights can be gained by looking at various details of the pressure field. Figure 18 shows the pressure contours at $t=7.6$ ms on a $yz$ plane at $x=0$, focusing on the pressure around the bubble and considering the three momentum balance modeling frameworks shown in Figure 17(b). In this snapshot, the bubble size and position is almost identical for all frameworks. Nonetheless, different pressure distributions and magnitudes are observed, with the simulation without momentum balance corrections seemingly predicting a lower pressure in the liquid overall. This is corroborated in Figure 19, which shows the pressure extracted along the diagonal of an $xy$ plane located at $z=0.00625$ mm, i.e., from $\mathbf{x}_{i}=(0,0,0.00625)$ mm to $\mathbf{x}_{f}=(4,4,0.00625)$ mm, at $t=7.6$ ms. Under the assumption that the bubble and its motion have negligible influence on the pressure at the considered depth, the hydrostatic pressure given the liquid column is 118.772 Pa. Indeed, the omission of the momentum balance corrections induces a lower pressure, while the addition of $\mathbf{f}_{NC}$, $\mathbf{f}_{\dot{m}^{\prime\prime}}$ and the Stefan flow shift improves the agreement.
Figure 18: Pressure field on the $yz$ plane with $x=0$ at $t=7.6$ ms of the rising ethanol bubble obtained with a grid spacing of 12.5 $\mu$m. The results without momentum balance corrections (left), only with the body forces $\mathbf{f}_{NC}$ and $\mathbf{f}_{\dot{m}^{\prime\prime}}$ (center), and with all corrections (right) are shown. The interface location is visualized by the iso-contour with $C=0.5$.
@@
However, further inspection reveals that the pressure field is affected by considerable oscillations if no momentum balance corrections are implemented, similar to Figure 6(c). Figure 20 shows the average pressure inside the bubble (i.e., from all pure gas cells), the average pressure around the bubble (i.e., from all pure liquid cells directly neighboring the interface) and the average pressure jump calculated by subtracting the two previous quantities. After an initial transient due to the relaxation of the initial conditions, the pressure inside and around the bubble obtained with all corrections implemented behaves smoothly over time. In contrast, frequent pressure oscillations are observed if the momentum balance is affected by the numerical approach. Here, the addition of only the body forces $\mathbf{f}_{NC}$ and $\mathbf{f}_{\dot{m}^{\prime\prime}}$ helps reducing the pressure oscillations after the initial transient although nearly complete mitigation is only possible once the Stefan flow shift is considered. It can be observed that at $t=7.6$ ms, the solution without corrections is predicting a smaller pressure than the trend, i.e., the observations from Figures 18 and 19. Moreover, the imbalance oscillations translate into pressure jump oscillations over time. During the initial transient, the magnitudes of the oscillations obtained without corrections and by only adding the body forces are very similar. This shows that the addition of the Stefan flow shift is responsible for improving the behavior of the pressure field during the adjustment phase too. Notably, taking proper care of the momentum balance is also reflected in the relatively smooth evolution of the average pressure jump.
@@
Lastly, it is worth noting that the trends of the pressure field quantities reported in Figure 20 are very similar during the early stages despite the oscillations introduced by the formulation without momentum balance corrections. This is more aligned with the expectation that the corrections are, in terms of magnitude, negligible for this problem. However, different rise velocities are reported. Since the interface dynamics strongly depend on the pressure jump between liquid and gas, it is expected that the bubble shape evolves differently if the solution of the momentum equation is according to the corrected pressure field. This results in a different balance between buoyancy and drag; thus, a different rise velocity is observed.
6 Assessment of the Viscous Stress Jump with Evaporation
A two-dimensional film boiling configuration is considered as an additional benchmark [61, 62, 17]. Another test fluid with the properties of fluid B shown in Table 3 is used. This highly viscous fluid causes a momentum imbalance across the interface due to the regularization approach of the viscous term assumed in Section 2.3, which is a more challenging task to address in this kind of multiphase solvers. As discussed below, the untreated discretization of the viscous term modifies the flow dynamics due to the sharp one-fluid formulation.
Following Rayleigh-Taylor instability theory, the vapor film is initialized with a perturbation of wavelength $\lambda_{d}$ to trigger the most unstable mode. The evaporation of the liquid increases the vapor volume and the dynamics are dominated by the combined effect of the momentum balance (e.g., surface tension) and buoyancy, forming a bubble that rises due to gravity (i.e., $\mathbf{g}=(0,-9.81)$ m/s2). The computational domain is a box of size $[0,\lambda_{d}/2]\times[0,\lambda_{d}]$ with symmetric boundaries at $x=0$ and $x=\lambda_{d}/2$, a wall with no-slip boundary conditions at $y=0$ and outflow boundary conditions at $y=\lambda_{d}$. For the pressure solver, homogeneous Neumann boundary conditions are imposed at the bottom and top boundaries with a reference pressure of 0 Pa at $\mathbf{x}_{p}=(0,\lambda_{d})$. For this problem, PLIC is used since no major differences are observed in the curvature computation and interface evolution with respect to PPIC; thus, favoring computational savings. The liquid remains at $T_{sat}$ and the wall temperature is fixed at $T_{wall}=T_{sat}+5$ K. The vapor film, which sits on top of the wall, is initialized with a linear temperature profile, $T_{0}$, between the initial interface location, $y_{0}$, and the wall. $\lambda_{d}$, $T_{0}$ and $y_{0}$ are given, respectively, by Eqs. (44), (45) and (46) (see [17] for more details). For the given fluid properties, $\lambda_{d}=78.6844$ mm.